OpenStudy (anonymous):

Variables on both sides. Solve the equation. 7-n=n-4-1

OpenStudy (kl0723):

move variables to the same side to make them easy to work with

OpenStudy (ahsome):

First, move the \(-4\) and the \(-1\) onto the LHS Before: \(7-n=n-4-1\) Now: \(7+4+1-n=n\) Simplify: \(12-n=n\) Move the \(n\) to the RHS \(12=n+n\) Simplify \(12=2n\) Solve: \(n=\frac{12}{2}\) \(n=6\)

OpenStudy (kl0723):

-n-n=-4-1-7

OpenStudy (ahsome):

Do you understand, @Jamiehughes?